Home‎ > ‎Crew 133‎ > ‎

111-CommandersReport

Commander Report
01/11/2014
Prepared by Paula Crock
 
We had another quiet in-Hab day today to keep up with the cleaning and work on projects, assembling
data and planning for the next round of experimentation.  In the morning Gartrelle and Diaz-Rubin
cleaned the EVA suit room while Howell and Komorowski brushed the mud off the EVA suits, gaiters and
boots and Jessup cleaned up the EVA suits and helmets.  I was floater, helping out as needed where
needed.  With teamwork the chores were finished before noon and we kicked back upstairs with a
beverage of choice (Gatorade has become popular lately) only to be dismayed by our internet
bandwidth – at breakfast we were at nearly 100% but now, just a few hours later and with all of us
busy doing chores, we were down to 12%!  As near as we can figure, when we brought in the
observatory computer late yesterday it got busy with a couple dozen windows updates, which chewed up
all of our bandwidth.  We’re still trying to sort it out, but we’re likely to go over our limit
today with our report submissions which means our communication with Earth will be severely limited
tomorrow.   Life on the frontier sometimes requires sacrifice; I suppose that’s part of the reason
why most everyone is back there and only a few of us are out here. We had another quiet in-Hab day
today to keep up with the cleaning and work on projects, assembling data and planning for the next
round of experimentation.  In the morning Gartrelle and Diaz-Rubin cleaned the EVA suit room while
Howell and Komorowski brushed the mud off the EVA suits, gaiters and boots and Jessup cleaned up the
EVA suits and helmets.  I was floater, helping out as needed where needed.  With teamwork the chores
were finished before noon and we kicked back upstairs with a beverage of choice (Gatorade has become
popular lately) only to be dismayed by our internet bandwidth – at breakfast we were at nearly 100%
but now, just a few hours later and with all of us busy doing chores, we were down to 12%!  As near
as we can figure, when we brought in the observatory computer late yesterday it got busy with a
couple dozen windows updates, which chewed up all of our bandwidth.  We’re still trying to sort it
out, but we’re likely to go over our limit today with our report submissions which means our
communication with Earth will be severely limited tomorrow.   Life on the frontier sometimes
requires sacrifice; I suppose that’s part of the reason why most everyone is back there and only a
few of us are out here.

--------------------------

CommanderCheckIn Report
01/11/2014
Prepared by Paula Crock

Crew Physical Health:
All are well.
  
Eva Depart Time:
N/A
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Hab Maint:
Attempting to track down source of sewer smell in Hab
Unknown internet bandwidth loss
Rep Sched:
Commander Check-In
Engineering Report
HSO Report
Science Report (GreenHab)
Future EVA Plan
Daily Photos
Journalist Report
Commander Report
Tomorrow Plan:
EVA #5 for Jessup and Crock projects
 
Inventory:
N/A
Support Req:
Still waiting on info on lights/ballast for lab lights
Miscelleneous:
N/A